7.48. Nitrogen narcosis begins to affect scuba divers at a depth of about 66 ft under water, when the total surround ing pressure is 4.0 atm. What is the solubility of N2 in blood at this depth, assuming that blood dissolves N2 like water?
Answer #1

Total pressure =4.0 \ atm

Mole fraction of nitrogen =0.78

Partial pressure of nitrogen =0.78 \times 4.0 \ atm=3.12 \ atm

The Henry's constant K= 6.8 \times 10^{-4} \ M/atm

Solubility of nitrogen in blood is S= KP= 6.8 \times 10^{-4} \ M/atm \times 3.12 \ atm=0.0021 \: M